OpenStudy (anonymous):

Amy's scores on the first four of five 100-point tests were 85, 89, 90, and 81. What score must she receive on the fifth test to have an average of at least 87 points for all the tests?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

345+x/5=87 345+x=87*5 345+x=435 x+90

OpenStudy (anonymous):

sorry x+90

OpenStudy (anonymous):

(85+89+90+81+x)/5>87 (345+x)>435 x>435-345 x>90